What is the molality of a solution containing 12.9 g Na2SO4 dissolved in 2980.2 g of water?
1
Expert's answer
2017-05-05T06:58:00-0400
Solution:
The molality of a solution – is the number of moles of solute in 1 kg of solvent (denoted mol/kg).
First we have to calculate the molar mass of Na2SO4 (M):
M = 23.0*2 + 32.1 + 16.0*4 = 142.1.

Therefore 12.9 g of it contain 12.9 / 142.1 = 0.09 mol.

So we have 0.09 mol of Na2SO4 in 2.9802 kg of water.
Than 1 kg of water contains 0.09mol / 2.9802kg = 0.03 mol/kg


Answer:
The molality of a solution is 0.03 mol/kg.